A force of 588 N acts on a box to move it 4m in 4o seconds.calculate the power.

Grade:

2 Answers

rajula vineet reddy
12 years ago

power=work done/timei.e power=F*S/T,power=588*4*40=94080watts

cute leo
12 years ago

But in my book answer is  58.8 watt.plz do correctly i have to submmit my work in this week